# VL-LN Bench: Towards Long-horizon Goal-oriented Navigation with Active Dialogs

### v1

@@ Abstract
In most existing embodied navigation tasks, instructions are well-defined and unambiguous, such as instruction following and object searching. Under this idealized setting, agents are required solely to produce effective navigation outputs conditioned on vision and language inputs. However, real-world navigation instructions are often vague and ambiguous, requiring the agent to resolve uncertainty and infer user intent through active dialog. To address this gap, we propose Interactive Instance Object Navigation (IION), a task that requires agents not only to generate navigation actions but also to produce language outputs via active dialog, thereby aligning more closely with practical settings. IION extends Instance Object Navigation (ION) by allowing agents to freely consult an oracle in natural language while navigating. Building on this task, we present the Vision Language–Language Navigation (VL-LN) benchmark, which provides a large-scale, automatically generated dataset and a comprehensive evaluation protocol for training and assessing dialog-enabled navigation models. VL-LN comprises over 41k long-horizon dialog-augmented trajectories for training and an automatic evaluation protocol with an oracle capable of responding to agent queries. Using this benchmark, we train a navigation model equipped with dialog capabilities and show that it achieves significant improvements over the baselines. Extensive experiments and analyses further demonstrate the effectiveness and reliability of VL-LN for advancing research on dialog-enabled embodied navigation. Code and dataset can be found at https://0309hws.github.io/VL-LN.github.io/.
Figure 1: A case for the IION task. The oracle (top left) first gives a simple goal-oriented navigation instruction (“Search for the chair.”). The agent has to locate a specific instance of the given category (chair). The agent can ask three types of questions—attribute, route, and disambiguation—to progressively resolve ambiguity and locate the target (instance). The full description in the bottom right is the instruction given to the agent in the ION task, which can locate the specific chair in this environment.
@@ I Introduction
A practical navigation agent must handle vague tasks by both planning effectively and resolving ambiguities. Active dialog offers a natural solution, allowing the agent to clarify underspecified instructions and obtain cues for efficient navigation. To study this ability, we propose Interactive Instance Object Navigation (IION; as shown in Fig. 1), which extends Instance Object Navigation (ION) [ION]. In IION, the agent receives a basic ObjectNav [objectnav] instruction (e.g.,“Search for the <category>”), which is insufficient to uniquely identify the target instance, and must consult an oracle via dialog. Since IION inherently poses exploration and disambiguation challenges, it provides an ideal testbed for examining the role of dialog in navigation.
@@
To effectively investigate and address the proposed IION task, this paper aims to build an Agent–Oracle interaction benchmark for dialog-enabled navigation agent training and evaluation. Prior work [thomason2020vision] collected human–human dialogs in navigation to evaluate whether agents can understand and follow instructions. However, such efforts do not assess an agent’s ability to proactively ask targeted questions. More recent approaches attempt to enable agents to generate questions, but they either limit the task to small-scale, room-level settings [gao2022dialfred] or focus narrowly on instance descriptions with limited support for long-horizon exploration [taioli2024collaborative]. Moreover, existing methods generally lack large-scale training datasets, which constrains the development of agents capable of both active exploration and informative questioning.
@@
In contrast to prior work, we propose the Vision Language–Language Navigation (VL-LN) benchmark, which provides both a large-scale, automatically generated training dataset and a comprehensive evaluation protocol for developing and assessing dialog-enabled navigation agents in long-horizon settings. The automated data collecting pipeline comprises three steps: (1) aggregating room-level region and instance attributes from MMScan [mmscan] into unified house-level annotations; (2) pairing target instances with feasible initial positions to instantiate episodes; and (3) collecting dialog-rich trajectories using a frontier-based navigator and a scripted oracle that answers questions to support exploration and disambiguation. Through this process, we curate the first large-scale training dataset of $\sim 41$K dialog-augmented trajectories for IION. The VL-LN benchmark further incorporates an evaluation protocol with the scripted oracle, enabling the assessment of agents’ dialog generation capabilities without requiring human intervention. Together, the dataset and evaluation protocol establish a unified benchmark for training and evaluating agents’ abilities in both language-based querying and navigation (LN) across IION and ION.
@@
We evaluate representative baselines on our benchmark and analyze emerging challenges. ION performance remains substantially lower than ObjectNav, highlighting two key difficulties: efficient exploration and instance-level disambiguation. Training on our collected IION data yields a language-enabled agent that achieves state-of-the-art results on both IION and ION, demonstrating the benefit of proactive querying. Nevertheless, the gains are not sufficient to bridge the disparity. A detailed error analysis reveals that the main bottleneck lies in image–attribute alignment, with 73% of failures caused by missing or misidentifying the target under detailed attributes. Moreover, in IION, the agent’s questioning is less efficient than humans’. Closing the gap needs better grounding, planning, and reasoning so the agent can identify the core ambiguity during exploration and pose maximally informative questions that efficiently reduce the candidate set. Our primary contributions are three-fold:
@@
  • An automatic pipeline for generating house-level, long-horizon IION trajectories.
@@
  • VL-LN, the first IION benchmark that provides a large-scale dialog-augmented dataset and online evaluation for both navigation and querying.
@@
  • Extensive experiments showing that active dialog improves agent performance on ION and IION, achieving the best results while revealing remaining challenges.
@@ II-A Goal-oriented navigation
Goal-oriented navigation requires an agent to find a specified goal in an unknown environment [ieong2025multimodal]. In the text-guided setting, it is commonly divided into Object-goal Navigation (ObjectNav) [anderson2018evaluation, objectnav] and Instance Object Navigation (ION) [ION]. Most prior work focuses on ObjectNav, where the instruction is unambiguous because any instance of the target category suffices. These approaches fall into two groups: training-based [campari2020exploiting, yadav2023offline, ye2021auxiliary, xie2023implicit, yin2025unigoal] and zero-shot [yokoyama2024vlfm]. Training-based methods typically use Imitation Learning [zare2024survey] or Reinforcement Learning [kaelbling1996reinforcement] to learn the observation-to-action mapping [ye2021auxiliary, xie2023implicit, yin2025unigoal], and developing vision–language aligned embeddings [campari2020exploiting, yadav2023offline] either in a single end-to-end model [yadav2023offline, ye2021auxiliary, xie2023implicit] or a modular pipeline [chen2023not, cai2024mo]. Zero-shot methods instead combine classical frontier-based exploration [yamauchi1997frontier] with the priors provided by large language [yu2023l3mvn, cai2024bridging] or vision–language models [majumdar2022zson, yokoyama2024vlfm] to score frontiers and guide exploration without task-specific training.
ION takes a step further by requiring the agent to locate a specific instance rather than any category member, which reflects another facet of real-world needs, where users often care about particular or personalized targets. [ION, barsellotti2024personalized]. Recent instance object navigation benchmarks [ION, barsellotti2024personalized] show that disambiguation is a central challenge, and that learned exploration policies often struggle to backtrack and recover after moving in the wrong direction. To address this problem, ION [ION] builds an instance-level graph whose nodes encode color, material, and location features, and PSL [sun2024prioritized] proposes Prioritized Semantic Learning for ION, which enhances semantic understanding to better identify the target. Despite these advances, real environments often contain many look-alike instances, and textual descriptions alone frequently lack the specificity needed to differentiate them [taioli2024collaborative]. Unlike prior work that treats the agent as a passive recipient of instance-level information, we enable the agent to proactively propose dialogs to refine the task specification and obtain targeted guidance, thus improving exploration efficiency.
@@ III VL-LN Benchmark
This section first defines the Interactive Instance Object Navigation (IION) task. We then present the VL-LN Benchmark, detailing the agent–oracle interaction in dialog-augmented trajectory generation and evaluation, the dataset construction for IION and dataset statistics, and the metrics employed to assess dialog quality.
@@ III-A Task Definition
The Interactive Instance Object Navigation (IION) challenges a dialog-enabled embodied agent to locate a specific instance in an unfamiliar environment. It involves two active roles: an agent and an oracle. For each episode, the agent is randomly placed in an unknown environment [szot2021habitat] and given an ambiguous instruction providing only the target category (e.g., “Search for the chair”). At each step, the agent receives a visual observation $o_{t}$ and odometry $l_{t}$ from the environment, and chooses to either move from $a_{t}\in\mathcal{A}$, where
@@
or query the oracle for guidance via free-form, open-ended natural-language interactions. The oracle is assumed to know all the information about the environment, such as: (1) the detailed attributes and location of the target instance, and (2) the global structure of the environment. The objective of the agent in IION is to locate the specified instance with minimal steps under a limited number of interactions.
TABLE I: Comparison with existing interactive instance object navigation benchmarks.
@@ III-B Agent–Oracle Interaction
In IION, the agent and oracle interact through template-free, open-ended natural-language dialogs. The agent may ask any question, and the oracle responds based on privileged access to environment information. To illustrate the scope of possible queries, consider the task of retrieving a friend’s computer in a large house, given only the instruction “Search for my computer.” To act efficiently, the agent may first request Attribute information to resolve ambiguity among same-category items (e.g., “What color is it?”). During exploration, the agent may then seek Route guidance to avoid blind searching (e.g., “Where should I go?”). Finally, upon encountering a candidate, the agent may ask a Disambiguation question to confirm correctness (e.g., “Is this the right computer?”). A positive confirmation therefore indicates successful task completion.
@@ III-C2 Episodes Generation
Each episode is defined by three core elements: an initial agent pose, a navigation instruction, and a set of target-instance viewpoints. The initial agent poses include starting poses from R2R-CE (72 scenes) [vln-ce] and our manually annotated 18 scenes. For the self-annotated scenes, we randomly sample navigable points, snap the agent to the selected location, and accept it as an initial pose only after verifying that the location is valid (on the navmesh, collision-free, and within scene bounds). For every instance, we provide two instruction variants: a partial, category-only instruction mentioned above and a full description that uniquely identifies the target among all instances in the scene (e.g., “Locate the deep grey chair with black backrest, standing upright on a wooden floor near a computer and a tv in the bedroom.”). To create the full description, we leverage the house-level dictionaries and the spatial-relation graph to select discriminative attributes and relations for the instance, and then prompt GPT-4o [gpt4o] to generate a natural-language instruction. The full instruction supports a non-interactive instance object navigation setting (i.e. ION), while the partial instruction is used in IION. We generate a set of viewpoints for each instance, which serve as expected “Stop” locations. Based on empirically reasonable viewing distances, we expand each instance’s 3D bounding box by $0.6\,\mathrm{m}$ in all directions and mark all navigable points within the expanded region as instance viewpoints. The episode is considered to be successful if the agent stops within 0.25 $m$ of any viewpoint of the instance. Combining an initial agent pose with the instance’s instruction and its viewpoints defines an episode.
@@ III-D Statistics
Our dataset covers 112 object categories that can serve as navigation targets (excluding structural elements such as walls and ceilings), comprising 20476 object instances across 90 annotated MP3D scenes. For each scene, we sample a set of navigable start poses (mean 42 per scene). Pairing these starts with the annotated instances yields 333,319 episodes. Following the VLN-CE partitioning [vln-ce], we allocate 61/15/14 scenes to train/val/test, resulting in 246,433 training episodes, 86,386 validation episodes, and 500 test episodes. Shown in Tab. I, we collect 41,891 trajectories spanning all scenes and categories of IION, each coupling navigation with question–answer interactions. Fig. 3 summarizes the trajectory statistics. In addition, we collect 5,087 ION and 23,774 ObjectNav trajectories generated by an FBE agent for the experiment.
@@ IV Experiments
In this section, we evaluate IION and ION using both zero-shot and trained baselines, analyzing why instance navigation is inherently difficult and how dialog contributes to reducing exploration failures and disambiguation errors, while highlighting remaining challenges in IION and providing some interesting findings.
@@ IV-A Experimental Setup
Evaluated task In addition to IION, we benchmark Instance Object Navigation (ION) without dialog, since most prior methods lack dialog capabilities. To enable the agent to identify the target instance, we provide the full instruction in the ION setting, which is a complete and unambiguous description that uniquely specifies the target.
Baselines We evaluate five baselines: two zero-shot and three training-based methods. The zero-shot baselines are (i) a greedy frontier-based exploration (FBE) agent that repeatedly selects the nearest frontier and uses an open-vocabulary detector built on Grounded SAM 2 [GroundingSam2] to detect the target instance, and (ii) VLFM [yokoyama2024vlfm], used as the released version. The learning-based baselines (VLLN-O, VLLN-I, VLLN-D) are initialized from Qwen2.5-VL-7B-Instruct and trained following the InternVLA-N1 [internvla-n1] procedure but with different data mixtures. All three include the VLN data from InternVLA-N1. VLLN-O additionally uses our ObjectNav data (23,774 trajectories), VLLN-I and VLLN-D further incorporate the filtered ION data (11,661 trajectories): VLLN-I trained without dialog, whereas VLLN-D trained with dialogs.

### v3

@@ Abstract
In most existing embodied navigation tasks, instructions are well-defined and unambiguous, such as instruction following and object searching. Under this idealized setting, agents are required solely to produce effective navigation outputs conditioned on vision and language inputs. However, real-world navigation instructions are often vague and ambiguous, requiring the agent to resolve uncertainty and infer user intent through active dialog. To address this gap, we propose Interactive Instance Goal Navigation (IIGN), a task that requires agents not only to generate navigation actions but also to produce language outputs via active dialog, thereby aligning more closely with practical settings. IIGN extends Instance Goal Navigation (IGN) by allowing agents to freely consult an oracle in natural language while navigating. Building on this task, we present the Vision Language–Language Navigation (VL-LN) benchmark, which provides a large-scale, automatically generated dataset and a comprehensive evaluation protocol for training and assessing dialog-enabled navigation models. VL-LN comprises over 41k long-horizon dialog-augmented trajectories for training and an automatic evaluation protocol with an oracle capable of responding to agent queries. Using this benchmark, we train a navigation model equipped with dialog capabilities and show that it achieves significant improvements over the baselines. Extensive experiments and analyses further demonstrate the effectiveness and reliability of VL-LN for advancing research on dialog-enabled embodied navigation. Code and dataset can be found at https://0309hws.github.io/VL-LN.github.io/.
Figure 1: A case for the IIGN task. The oracle (top left) first gives a simple goal-oriented navigation instruction (“Search for the chair.”). The agent has to locate a specific instance of the given category (chair). The agent can ask three types of questions—attribute, route, and disambiguation—to progressively resolve ambiguity and locate the target (instance). The full description in the bottom right is the instruction given to the agent in the IGN task, which can locate the specific chair in this environment.
@@ I Introduction
A practical navigation agent must handle vague tasks by both planning effectively and resolving ambiguities. Active dialog offers a natural solution, allowing the agent to clarify underspecified instructions and obtain cues for efficient navigation. To study this ability, we propose Interactive Instance Goal Navigation (IIGN; as shown in Fig. 1), which extends Instance-level Object Navigation (ION, named as IGN in our paper) [ION]. In IIGN, the agent receives a basic ObjectNav [objectnav] instruction (e.g.,“Search for the <category>”), which is insufficient to uniquely identify the target instance, and must consult an oracle via dialog. Since IIGN inherently poses exploration and disambiguation challenges, it provides an ideal testbed for examining the role of dialog in navigation.
@@
To effectively investigate and address the proposed IIGN task, this paper aims to build an Agent–Oracle interaction benchmark for dialog-enabled navigation agent training and evaluation. Prior work [thomason2020vision] collected human–human dialogs in navigation to evaluate whether agents can understand and follow instructions. However, such efforts do not assess an agent’s ability to proactively ask targeted questions. More recent approaches attempt to enable agents to generate questions, but they either limit the task to small-scale, room-level settings [gao2022dialfred] or focus narrowly on instance descriptions with limited support for long-horizon exploration [taioli2024collaborative]. Moreover, existing methods generally lack large-scale training datasets, which constrains the development of agents capable of both active exploration and informative questioning.
@@
In contrast to prior work, we propose the Vision Language–Language Navigation (VL-LN) benchmark, which provides both a large-scale, automatically generated training dataset and a comprehensive evaluation protocol for developing and assessing dialog-enabled navigation agents in long-horizon settings. The automated data collecting pipeline comprises three steps: (1) aggregating room-level region and instance attributes from MMScan [mmscan] into unified house-level annotations; (2) pairing target instances with feasible initial positions to instantiate episodes; and (3) collecting dialog-rich trajectories using a frontier-based navigator and a scripted oracle that answers questions to support exploration and disambiguation. Through this process, we curate the first large-scale training dataset of $\sim 41$K dialog-augmented trajectories for IIGN. The VL-LN benchmark further incorporates an evaluation protocol with the scripted oracle, enabling the assessment of agents’ dialog generation capabilities without requiring human intervention. Together, the dataset and evaluation protocol establish a unified benchmark for training and evaluating agents’ abilities in both language-based querying and navigation (LN) across IIGN and IGN.
@@
We evaluate representative baselines on our benchmark and analyze emerging challenges. IGN performance remains substantially lower than ObjectNav, highlighting two key difficulties: efficient exploration and instance-level disambiguation. Training on our collected IIGN data yields a language-enabled agent that achieves state-of-the-art results on both IIGN and IGN, demonstrating the benefit of proactive querying. Nevertheless, the gains are not sufficient to bridge the disparity. A detailed error analysis reveals that the main bottleneck lies in image–attribute alignment, with 73% of failures caused by missing or misidentifying the target under detailed attributes. Moreover, in IIGN, the agent’s questioning is less efficient than humans’. Closing the gap needs better grounding, planning, and reasoning so the agent can identify the core ambiguity during exploration and pose maximally informative questions that efficiently reduce the candidate set. Our primary contributions are three-fold:
@@
  • An automatic pipeline for generating house-level, long-horizon IIGN trajectories.
@@
  • VL-LN, the first IIGN benchmark that provides a large-scale dialog-augmented dataset and online evaluation for both navigation and querying.
@@
  • Extensive experiments showing that active dialog improves agent performance on IGN and IIGN, achieving the best results while revealing remaining challenges.
@@ II-A Goal-oriented navigation
Goal-oriented navigation requires an agent to find a specified goal in an unknown environment [ieong2025multimodal]. In the text-guided setting, it is commonly divided into Object-goal Navigation (ObjectNav) [anderson2018evaluation, objectnav] and Instance-level Object Navigation (ION) [ION]. Most prior work focuses on ObjectNav, where the instruction is unambiguous because any instance of the target category suffices. These approaches fall into two groups: training-based [campari2020exploiting, yadav2023offline, ye2021auxiliary, xie2023implicit, yin2025unigoal] and zero-shot [yokoyama2024vlfm]. Training-based methods typically use Imitation Learning [zare2024survey] or Reinforcement Learning [kaelbling1996reinforcement] to learn the observation-to-action mapping [ye2021auxiliary, xie2023implicit, yin2025unigoal], and developing vision–language aligned embeddings [campari2020exploiting, yadav2023offline] either in a single end-to-end model [yadav2023offline, ye2021auxiliary, xie2023implicit] or a modular pipeline [chen2023not, cai2024mo]. Zero-shot methods instead combine classical frontier-based exploration [yamauchi1997frontier] with the priors provided by large language [yu2023l3mvn, cai2024bridging] or vision–language models [majumdar2022zson, yokoyama2024vlfm] to score frontiers and guide exploration without task-specific training.
IGN takes a step further by requiring the agent to locate a specific instance rather than any category member, which reflects another facet of real-world needs, where users often care about particular or personalized targets. [ION, barsellotti2024personalized]. Recent instance goal navigation benchmarks [ION, barsellotti2024personalized] show that disambiguation is a central challenge, and that learned exploration policies often struggle to backtrack and recover after moving in the wrong direction. To address this problem, IGN [ION] builds an instance-level graph whose nodes encode color, material, and location features, and PSL [sun2024prioritized] proposes Prioritized Semantic Learning for IGN, which enhances semantic understanding to better identify the target. Despite these advances, real environments often contain many look-alike instances, and textual descriptions alone frequently lack the specificity needed to differentiate them [taioli2024collaborative]. Unlike prior work that treats the agent as a passive recipient of instance-level information, we enable the agent to proactively propose dialogs to refine the task specification and obtain targeted guidance, thus improving exploration efficiency.
@@ III VL-LN Benchmark
This section first defines the Interactive Instance Goal Navigation (IIGN) task. We then present the VL-LN Benchmark, detailing the agent–oracle interaction in dialog-augmented trajectory generation and evaluation, the dataset construction for IIGN and dataset statistics, and the metrics employed to assess dialog quality.
@@ III-A Task Definition
The Interactive Instance Goal Navigation (IIGN) challenges a dialog-enabled embodied agent to locate a specific instance in an unfamiliar environment. It involves two active roles: an agent and an oracle. For each episode, the agent is randomly placed in an unknown environment [szot2021habitat] and given an ambiguous instruction providing only the target category (e.g., “Search for the chair”). At each step, the agent receives a visual observation $o_{t}$ and odometry $l_{t}$ from the environment, and chooses to either move from $a_{t}\in\mathcal{A}$, where
@@
or query the oracle for guidance via free-form, open-ended natural-language interactions. The oracle is assumed to know all the information about the environment, such as: (1) the detailed attributes and location of the target instance, and (2) the global structure of the environment. The objective of the agent in IIGN is to locate the specified instance with minimal steps under a limited number of interactions.
TABLE I: Comparison with existing interactive instance goal navigation benchmarks.
@@ III-B Agent–Oracle Interaction
In IIGN, the agent and oracle interact through template-free, open-ended natural-language dialogs. The agent may ask any question, and the oracle responds based on privileged access to environment information. To illustrate the scope of possible queries, consider the task of retrieving a friend’s computer in a large house, given only the instruction “Search for my computer.” To act efficiently, the agent may first request Attribute information to resolve ambiguity among same-category items (e.g., “What color is it?”). During exploration, the agent may then seek Route guidance to avoid blind searching (e.g., “Where should I go?”). Finally, upon encountering a candidate, the agent may ask a Disambiguation question to confirm correctness (e.g., “Is this the right computer?”). A positive confirmation therefore indicates successful task completion.
@@ III-C2 Episodes Generation
Each episode is defined by three core elements: an initial agent pose, a navigation instruction, and a set of target-instance viewpoints. The initial agent poses include starting poses from R2R-CE (72 scenes) [vln-ce] and our manually annotated 18 scenes. For the self-annotated scenes, we randomly sample navigable points, snap the agent to the selected location, and accept it as an initial pose only after verifying that the location is valid (on the navmesh, collision-free, and within scene bounds). For every instance, we provide two instruction variants: a partial, category-only instruction mentioned above and a full description that uniquely identifies the target among all instances in the scene (e.g., “Locate the deep grey chair with black backrest, standing upright on a wooden floor near a computer and a tv in the bedroom.”). To create the full description, we leverage the house-level dictionaries and the spatial-relation graph to select discriminative attributes and relations for the instance, and then prompt GPT-4o [gpt4o] to generate a natural-language instruction. The full instruction supports a non-interactive instance goal navigation setting (i.e. IGN), while the partial instruction is used in IIGN. We generate a set of viewpoints for each instance, which serve as expected “Stop” locations. Based on empirically reasonable viewing distances, we expand each instance’s 3D bounding box by $0.6\,\mathrm{m}$ in all directions and mark all navigable points within the expanded region as instance viewpoints. The episode is considered to be successful if the agent stops within 0.25 $m$ of any viewpoint of the instance. Combining an initial agent pose with the instance’s instruction and its viewpoints defines an episode.
@@ III-D Statistics
Our dataset covers 112 object categories that can serve as navigation targets (excluding structural elements such as walls and ceilings), comprising 20476 object instances across 90 annotated MP3D scenes. For each scene, we sample a set of navigable start poses (mean 42 per scene). Pairing these starts with the annotated instances yields 333,319 episodes. Following the VLN-CE partitioning [vln-ce], we allocate 61/15/14 scenes to train/val/test, resulting in 246,433 training episodes, 86,386 validation episodes, and 500 test episodes. Shown in Tab. I, we collect 41,891 trajectories spanning all scenes and categories of IIGN, each coupling navigation with question–answer interactions. Fig. 3 summarizes the trajectory statistics. In addition, we collect 5,087 IGN and 23,774 ObjectNav trajectories generated by an FBE agent for the experiment.
@@ IV Experiments
In this section, we evaluate IIGN and IGN using both zero-shot and trained baselines, analyzing why instance navigation is inherently difficult and how dialog contributes to reducing exploration failures and disambiguation errors, while highlighting remaining challenges in IIGN and providing some interesting findings.
@@ IV-A Experimental Setup
Evaluated task In addition to IIGN, we benchmark Instance goal Navigation (IGN) without dialog, since most prior methods lack dialog capabilities. To enable the agent to identify the target instance, we provide the full instruction in the IGN setting, which is a complete and unambiguous description that uniquely specifies the target.
Baselines We evaluate five baselines: two zero-shot and three training-based methods. The zero-shot baselines are (i) a greedy frontier-based exploration (FBE) agent that repeatedly selects the nearest frontier and uses an open-vocabulary detector built on Grounded SAM 2 [GroundingSam2] to detect the target instance, and (ii) VLFM [yokoyama2024vlfm], used as the released version. The learning-based baselines (VLLN-O, VLLN-I, VLLN-D) are initialized from Qwen2.5-VL-7B-Instruct and trained following the InternVLA-N1 [internvla-n1] procedure but with different data mixtures. All three include the VLN data from InternVLA-N1. VLLN-O additionally uses our ObjectNav data (23,774 trajectories), VLLN-I and VLLN-D further incorporate the filtered IGN data (11,661 trajectories): VLLN-I trained without dialog, whereas VLLN-D trained with dialogs.